OpenStudy (anonymous):

Sally purchased a dining room set for $1,780 using a 12-month deferred payment plan with an interest rate of 25.48%. She did not make any payments during the deferment period. What will Sally’s monthly payment be if she must pay off the dining room set within two years after the deferment period?

OpenStudy (tkhunny):

She should have made payments! Yikes. \[1780\cdot (1.2548)=PMT\frac{1−v^{12}}{i}\] where i = 0.2548/12 and v = 1/(1+i). Is this looking familiar?

OpenStudy (anonymous):

I have no idea how to even start this problem.

OpenStudy (tkhunny):

Start?! It's almost done. Accumulate interest for one year then calculate the monthly payment necessary to pay it off in year 2. Is that how you read the problem statement?

OpenStudy (anonymous):

$122.80 is this the answer?

OpenStudy (tkhunny):

$122.80*12 = $1,473.60 - No. That will not pay off the debt in year 2. Much more.

OpenStudy (anonymous):

$105.35 $122.80 $74.17 $95.44 Those are the options.

OpenStudy (tkhunny):

I was afraid of that. The answer is printed incorrectly. The correct answer is $212.80, just a transposition error away from choice #2. (Really, $212.81)

OpenStudy (anonymous):

Thank you so much for helping me :)